Show that if X and Y are sequences such that X and X + Y are convergent, then Y is convergent.
1
Expert's answer
2020-04-22T18:57:23-0400

Since X = {xn} n=1,2,.. is convergent< we obtain that there exists A = "\lim\limits_{n\to\infin} x_n". So, for every "\varepsilon">0 there exists such N0, that for every n>N0 the following is true: |xn - A|<"\frac{\varepsilon}{2}" . (1)


We obtain the same for X+Y = {xn + yn}, n = 1,2,.. and B = "\lim\limits_{n\to\infin} (x_n + y_n)" :

for every "\varepsilon > 0" there exists such N1, that for every n>N1 the following is true:

|xn + yn - B|<"\frac{\varepsilon}{2} (2)"


Let us now denote with N = max{N0, N1}. Let's show that Y is convergent with the following limit:

B-A = "\lim\limits_{n\to\infin} y_n" . For this, let's check the definition:

We will prove that for every "\varepsilon > 0" there exists such N2, that for every n>N2 the following is true:

|yn - (B-A)|<"\varepsilon" (the definition of the limit): (assuming n>N (1) and (2) are both true)


|yn - (B-A)| = |(xn - A) - (xn + yn - B)| "\leqslant" |xn - A| + |xn + yn - B| "\leqslant" "\frac{\varepsilon}{2} + \frac{\varepsilon}{2} = \varepsilon". So for n>N this is true.


Exactly what we needed to show.